A. Directions: Tossing four coins. Suppose four coins are tossed. Let Y be





the random variable representing the number of tails that occur. Complete





the table below by showing the values of random variable Y.





Let H = Heads





T = Tails

1
Expert's answer
2022-03-08T05:14:03-0500

The possible number of tails (Y) is: 0,1,2,3,4.

Computing the respective probabilities using the binomial distribution and assuming that the probability of a tail and the probability of a head are equal:

  1. "P(Y=0)=\left(\frac{1}{2}\right)^4=\frac{1}{16}"
  2. "P(Y=1)=C_4^1\left(\frac{1}{2}\right)^4=4\left(\frac{1}{2}\right)^4=\frac{1}{4}"
  3. "P(Y=2)=C_4^2\left(\frac{1}{2}\right)^4=\frac{3\cdot4}{2}\left(\frac{1}{2}\right)^4=\frac{3}{8}"
  4. "P(Y=3)=C_4^3\left(\frac{1}{2}\right)^4=\frac{1}{4}"
  5. "P(Y=4)=C_4^4\left(\frac{1}{2}\right)^4=\frac{1}{16}"


The Probability distribution of Y

"\\begin{matrix}Y&{0}&{1}&{2}&{3}&{4}\\\\p&{\\frac{1}{{16}}}&{\\frac{1}{{4}}}&{\\frac{3}{{8}}}&{\\frac{1}{{4}}}&{\\frac{1}{{16}}}\\end{matrix}\n\u200b"